# Holistic Optimization Framework for FPGA Accelerators

### v4

@@ Abstract.
High-Level Synthesis, NonLinear Programming, Compiler
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Hardware High-level and register-transfer level synthesisccs: Software and its engineering Compilers
@@ 1. Introduction
To address this complexity, High-Level Synthesis (HLS) tools and source-to-source compilers (Cong et al., 2011; Zhang et al., 2008; Xilinx, 0232; Intel, 2024; Siemens, 2023; Microchip, 2023; Xilinx, 2024; Huang et al., 2021; Ye et al., 2022; Lai et al., 2019; Yu et al., 2018; Li et al., 2020; Pouget et al., 2025b; Basalama and Cong, 2025) have emerged as critical enablers of FPGA adoption, allowing developers to write hardware-accelerated programs in high-level languages such as C++ and Python. These tools generate synthesizable hardware descriptions, leveraging pragmas (hardware directives) and code transformations to optimize execution. Despite their advantages, HLS-generated designs still require substantial manual tuning to achieve high QoR, as the compiler’s ability to optimize performance remains highly dependent on the structure and directives applied to the input code.
@@ 1.1. Challenges in FPGA Optimization
Scalability and Multi-SLR Utilization: FPGA architectures, especially those using Stacked Silicon Interconnect (SSI) technology, feature multiple Super Logic Regions (SLRs) to enhance scalability. However, most frameworks, including Merlin-based tools (e.g., AutoDSE, HARP, NLP-DSE) and Sisyphus, restrict designs to a single SLR, leading to underutilization of FPGA resources. While designing across multiple SLRs is possible, it significantly increases routing congestion, timing closure failures, and bitstream generation complexity. Without an SLR-aware optimization strategy, large-scale tasks face a high probability of bitstream generation failures or degraded performance due to excessive inter-SLR communication overhead.
@@ 1.2. Contributions and Prometheus Method
This work introduces Prometheus, a unified framework that automates FPGA design space exploration by integrating multiple optimizations into a single methodology. Unlike existing tools that optimize isolated aspects of FPGA acceleration, Prometheus holistically addresses code transformation, memory management, task concurrency, and hardware-aware scheduling. It leverages a hybrid execution model that balances shared memory reuse and dataflow streaming to maximize parallelism while minimizing memory overhead.
@@
The methodology follows a structured process: First, Prometheus performs affine code analysis to identify parallelism opportunities and dependencies. It then generates a dataflow graph and fuses tasks that share the same outputs to minimize unnecessary communication overhead. Next, Prometheus constructs a comprehensive design space that includes all key optimizations outlined in Section 2, such as task concurrency via dataflow, computation-communication overlap, and adaptive parallelism, while ensuring resource constraints are met for each SLR. The NLP-based design space exploration model, as described in Section 4, is then employed to identify the theoretically optimal set of transformations. Once the best configuration is identified, Prometheus automatically generates the optimized design. Finally, the system produces an OpenCL host code and compiles the FPGA bitstream with the HLS compiler.
@@
  • Unified FPGA optimization framework that jointly optimizes loop transformations, pragma insertion, task concurency and computation-communication overlap while considering interdependencies between computation and data movement.
@@
  • NLP-based design space exploration that automates the selection of loop tiling, scheduling, and memory strategies for globally theoretical optimal performance.
@@
  • End-to-end compilation and automation that generates optimized HLS-C++ code, OpenCL host code, and FPGA bitstreams with minimal manual intervention.
@@ 2. Background and Motivation
Pragma Insertion Limit.
@@
Task Concurrency Limit.
@@
Computation-Communication Overlap Limit.
@@
Hardware Aware Limit.
Hardware-feasible Limit. Limit. Limit. Limit.
@@ 2.1.3. Task Concurrency
For computational kernels such as 3mm (Listing 3), the dataflow paradigm allows the first two matrix multiplications to execute in parallel while the third begins as soon as its required inputs are produced. This overlapping of execution helps maximize throughput and minimize idle time for computing units. Additionally, dataflow optimizations facilitate task-level parallelism, allowing independent tasks to run concurrently across multiple compute resources, such as DSP blocks and BRAM, ensuring efficient utilization of available FPGA resources.
@@
Stream-HLS (Basalama and Cong, 2025) attempts to address this limitation by increasing the number of FIFOs connecting two tasks, assuming that all data resides on-chip. While this approach simplifies certain aspects of the design space, it is neither scalable nor generalizable for transferring data from off-chip to on-chip memory. By using $n$ FIFOs to transfer $n$ data elements in parallel between two tasks, the method significantly increases resource consumption without proportionally improving efficiency. The additional FIFOs complicate the design, potentially leading to routing congestion and excessive hardware overhead.
@@ 2.1.4. Shared Buffering
This approach involves preloading data buffers into on-chip memory, allowing multiple computations to access shared data without redundant transfers. By reducing memory access latency and improving bandwidth utilization, shared buffering facilitates high parallelism and optimizes overall performance. It is particularly beneficial in applications which are computation bound.
@@ 2.1.6. Data Packing and Padding
Modern FPGA architectures support high-bandwidth data transfers (up to 512-bit wide on AMD/Xilinx FPGAs). Data packing and padding optimize memory alignment, reducing the number of required memory cycles. For instance, transferring 216 floating-point values using a 256-bit width (8 floats per transfer) requires 27 cycles, making efficient packing crucial for minimizing overhead.
@@
For example, in the code presented in Listings 1 and 2, where $J=190$, transferring all elements along the second dimension (iterated by loop $j$) of array $B$ onto the chip is constrained by the available memory bandwidth. Without padding, the maximum transfer rate is $64$ bits per cycle, as $190\times 32$ is divisible by $64$ but not by $128$. However, by introducing padding with $P=2$ and adjusting $J$ to $192$ ($190+2$), the data alignment enables a significantly higher transfer rate of $512$ bits per cycle, as $192\times 32$ is now perfectly divisible by $512$. This optimization maximizes memory bandwidth utilization, reducing transfer latency and improving overall data throughput.
@@
Padding can also be used for computation, if we take a similar method than Sisyphus (Pouget et al., 2025b) which tiled and unroll the intra-tile loops which correspond to the transformation from Listing 1 to Listing 2. We will have an unroll factor which correspond to $I1\times J1\times K1$ but we want to have $I1$ divide $I$, $J1$ divide $J$ and $K1$ divide K because we do not want to use extra ressource to only compute a partial tile which would correspond to the execution of the last tile which is not complete. For example if we have a trip count of $190$ and we use an unroll factor of $8$ then we will execute $184\times 8$ a full tile and then 6 iterations for the partial tile.
But we restrict the unroll factor to be divisible of the trip count we restrict considerably the space of possibilities. So the solution is to pad which allow to considerably increase the space of possibilities. Hence with a trip count of $190$ the possible unroll factor are include in $UF\in\{1,2,5,10,19,38,95,190\}$ but with padding of 2 (a trip count of 192) the space become $UF\in\{1,2,3,4,6,8,12,16,24,32,48,64,96,192\}$. And of course all pading can be considered. Padding enables finer control over unrolling and resource utilization, allowing for more flexible and efficient hardware implementation while optimizing computational parallelism.
@@ 4 C[i][j] += A[i][k] * B[k][j];
Listing 1: Baseline Implementation of Matrix Multiplication in C. This code depicts a naive triple-nested loop structure used for matrix-matrix multiplication, serving as the unoptimized reference for further transformations.
@@ 10 C[i0*I1 + i1][j0*J1 + j1] += A[i0*I1 + i1][k0*K1 + k1] * B[k0*K1 + k1][j0*J1 + j1];
Listing 2: Matrix Multiplication with Loop Tiling and Fully Unrolled Intra-Tile Computation. This implementation showcases a performance-optimized matrix multiplication using loop tiling to enhance data locality and fully unrolled intra-tile loops to expose fine-grained parallelism.
@@ 2.3. Challenges
A common oversimplification in many frameworks is the assumption that all data resides on-chip. While this simplifies memory access patterns in design space exploration, it often results in low Quality of Results (QoR) when deployed on real hardware. This assumption prevents frameworks from incorporating crucial tiling strategies, which are essential for optimizing data locality, reducing memory access overhead, and ensuring efficient off-chip communication. When off-chip memory management is treated as a separate optimization step, decisions made in earlier phases may restrict the effectiveness of later memory optimizations, leading to suboptimal performance and resource utilization.
The limitations of current HLS optimization frameworks that rely on shared buffering are exemplified through the 3mm kernel. It consists of two independent matrix multiplications, whose outputs serve as inputs for a final matrix multiplication.
As illustrated in Listing 3, each matrix multiplication consists of deeply nested loops that are well-suited to parallelization techniques such as loop unrolling, pipelining, and computation-communication overlap. However, conventional shared buffering strategies often fall short in effectively leveraging concurrency, primarily due to their limited exploitation of dataflow principles and insufficient overlap between computation and communication.
@@
Allo (Chen et al., 2024) aims to reduce development time and enhance design quality. However, it still requires significant manual intervention from an expert, limiting its usability for non-specialists.
@@ 15 G[i][j] += E[i][k] * F[k][j];} // S5
Listing 3: Reference Implementation of the 3mm Kernel from PolyBench. This listing shows the baseline version of the 3mm benchmark, comprising three matrix multiplication operations. The first two multiplications are independent and produce intermediate results that serve as inputs to the final matrix multiplication.
@@ 2.4. Overview of Prometheus
On-Chip Memory Total memory usage by buffered tiles (considering reuse and buffering) must be within BRAM capacity.
@@
Figure 1 illustrates the workflow of Prometheus. The process begins with an input C++ code, from which we extract an intermediate representation containing dependencies, trip counts, schedules, and other relevant information. This extraction is performed using PoCC (Pouchet, 2023), which provides all necessary details for analysis. Next, we construct the task-flow graph based on this extracted information and formulate the corresponding Non-Linear Programming (NLP) problem. Solving this NLP problem determines the theoretical optimal parameters for the design space. Once the NLP solver provides these parameters, we have all the required information to proceed. At this stage, we automatically generate the HLS-C++ code along with all the necessary files to produce the FPGA bitstream, ensuring an efficient and fully automated compilation process.
@@
The pseudo-code in Listing 4 demonstrates how Prometheus processes the 3mm kernel (Listing 3). Load and store operations manage data transfers to and from off-chip memory, while send and receive operations handle inter-task communication using FIFO.
Each taski in the pseudo-code corresponds to the fully unrolled computation of an intra-tile for statement Si in Listing 3. An example of such a task is illustrated in Listing 5.
@@
Prometheus enhances computation efficiency by fusing statements that produce the same outputs (e.g., in Listing 3), and it automatically formulates an NLP problem to determine the theoretically optimal parameters, such as loop schedules, array bit widths (e.g., 512 bits), tile sizes, reuse buffer sizes, transfer locations, and padding. The framework dynamically adjusts bit widths for arrays like $F$, $D$, and $G$ to achieve a better balance between parallelism and resource usage.
@@
Listing 4: Transformed and Fused 3mm Kernel Pseudocode Generated by Prometheus. This code highlights the result of Prometheus optimizations, including fused task generation, tiled loops, memory buffer management, and pipelined execution with dataflow pragmas.

### v5

@@ Abstract.
High-Level Synthesis, Non-Linear Programming, Compiler
copyright: ccjournal: TODAESdoi: 10.1145/3769307ccs: Hardware High-level and register-transfer level synthesisccs: Software and its engineering Compilers
@@ 1. Introduction
To address this complexity, High-Level Synthesis (HLS) tools (Xilinx, 0232; Intel, 2024; Siemens, 2023; Microchip, 2023) and source-to-source compilers e.g,  (Ye et al., 2022; Lai et al., 2019; Li et al., 2020; Pouget et al., 2025b; Basalama and Cong, 2025), have emerged as critical enablers of FPGA adoption, allowing developers to write hardware-accelerated programs in high-level languages such as C++ and Python. These tools generate synthesizable hardware descriptions, leveraging pragmas (hardware directives) and code transformations to optimize execution. Despite their advantages, HLS-generated designs still require substantial manual tuning to achieve high QoR, as the compiler’s ability to optimize performance remains highly dependent on the structure and directives applied to the input code.
@@ 1.1. Challenges in FPGA Optimization
Scalability and Multi-Super Logic Regions Utilization: FPGA architectures, especially those using Stacked Silicon Interconnect (SSI) technology, feature multiple Super Logic Regions (SLRs) to enhance scalability. However, most frameworks, including Merlin-based tools restrict designs to a single SLR, leading to underutilization of FPGA resources. While designing across multiple SLRs is possible, it significantly increases routing congestion, timing closure failures, and bitstream generation complexity. Without an SLR-aware optimization strategy, large-scale tasks face a high probability of bitstream generation failures or degraded performance due to excessive inter-SLR communication overhead.
@@ 1.2. Contributions and Prometheus Method
This work introduces Prometheus, a unified framework that automates FPGA design space exploration by integrating multiple optimizations into a single methodology specialized for affine programs (Lattner et al., 2021). Specifically, we operate on affine loop nests where statements can be distributed in multiple loops (i.e. loops are permutable (Pouchet et al., 2011)). Such programs encompass typical computational pattern in dense linear algebra, data mining, image processing, etc. as exemplified in the PolyBench benchmarking suite (Polybench, [n. d.]; Abella-González et al., 2021). Unlike existing tools that optimize isolated aspects of FPGA acceleration such as AutoDSE (Sohrabizadeh et al., 2021), HARP (Sohrabizadeh et al., 2023), NLP-DSE (Pouget et al., 2025a), Sisyphus (Pouget et al., 2025b) and Stream-HLS (Basalama and Cong, 2025), Prometheus holistically addresses code transformation, memory management, task concurrency, and hardware-aware scheduling. It leverages a hybrid execution model that balances shared memory reuse and dataflow streaming to maximize parallelism while minimizing memory overhead.
@@
The methodology follows a structured process: First, Prometheus performs affine code analysis to identify parallelism opportunities and dependencies. It then generates a dataflow graph and fuses tasks that share the same outputs to minimize unnecessary communication overhead. Next, Prometheus constructs a comprehensive design space that includes all key optimizations outlined in Section 2, such as task concurrency via dataflow, computation-communication overlap, and adaptive parallelism, while ensuring resource constraints are met for each SLR. The NLP-based design space exploration model, as described in Section 4, is then employed to identify the theoretically optimal set of transformations. Once the best configuration is identified, Prometheus automatically generates the optimized design. Finally, the system produces OpenCL host code and compiles the FPGA bitstream with the HLS compiler.
@@
  • Unified FPGA optimization framework that jointly optimizes loop transformations, pragma insertion, task concurrency and computation-communication overlap while considering interdependencies between computation and data movement.
@@
  • NLP-based design space exploration that automates the selection of loop tiling, scheduling, and memory strategies to achieve a globally optimal theoretical performance.
@@
  • End-to-end compilation and automation, generating optimized HLS-C++ code—that is, standard C++ source code annotated with AMD/Vitis-specific high-level synthesis (HLS) pragmas to guide hardware generation—alongside OpenCL host code and FPGA bitstreams, all with minimal manual intervention.
@@
Our framework, Prometheus, is open source and available at https://github.com/UCLA-VAST/Prometheus.
@@ 2. Background and Motivation
Pragma Insertion Limit
@@
Task Concurrency Limit
@@
Computation-Communication Overlap Limit
@@
Hardware Aware Limit
Hardware-feasible Limit Limit Limit Limit
@@ 2.1.3. Task Concurrency
For computational kernels such as 3mm (Listing 4), the dataflow paradigm allows the first two matrix multiplications to execute in parallel while the third begins as soon as its required inputs are produced. This overlapping of execution helps maximize throughput and minimize idle time for computing units. Additionally, dataflow optimizations facilitate task-level parallelism, allowing independent tasks to run concurrently across multiple compute resources, such as DSP blocks and BRAM, ensuring efficient utilization of available FPGA resources.
@@
Stream-HLS (Basalama and Cong, 2025) addressed this limitation by increasing the number of FIFOs connecting two tasks, assuming that all data resides on-chip. While this approach simplifies certain aspects of the design space, it is neither scalable nor generalizable for transferring data from off-chip to on-chip memory. By using $n$ FIFOs to transfer $n$ data elements in parallel between two tasks, the method significantly increases resource consumption without proportionally improving efficiency. The additional FIFOs complicate the design, potentially leading to routing congestion and excessive hardware overhead.
@@ 2.1.4. Shared Buffering
This approach involves preloading data buffers into on-chip memory, allowing multiple computations to access shared data without redundant transfers. By reducing memory access latency and improving bandwidth utilization, shared buffering facilitates high parallelism and optimizes overall performance. It is particularly beneficial in applications that are computation-bound.
@@ 2.1.6. Data Packing and Padding
Modern FPGA architectures support high-bandwidth data transfers (up to 512-bit wide on AMD/Xilinx FPGAs). Data packing and padding optimize memory alignment, reducing the number of required memory cycles. For instance, transferring 216 floating-point values using a 256-bit width (8 floats per transfer) requires 27 cycles – compared to 216 cycles without packing. This highlights the importance of efficient data packing to minimize overhead. However, to enable such packing, the transfer vector size must evenly divide the total array size.
@@
Figure 1. Illustration of Padding Strategy to Align Array Size for Efficient Data Packing and Memory Transfers on AMD FPGAs
1// Original loop without padding
2for (int i = 0; i < 190; i++){
3#pragma HLS unroll factor=UF
4// Possible UF values: 1, 2, 5, 10, 19, 38, 95, 190
5 A[i] += 1;
6}
7
8// Padded loop
9for (int i = 0; i < 190 + 2; i++){ // Padding of 2
10#pragma HLS unroll factor=UF
11// Possible UF values: 1, 2, 3, 4, 6, 8, 12, 16, 24, 32, 48, 64, 96, 192
12 A[i] += 1;
13}
Listing 1: Effect of Padding on the Space of Unroll Factors for Computation
Figure 1 illustrates the selection of the padding value $P$ for an array $A$ of size $N$. By leveraging data packing (up to 512 bits), multiple elements can be transferred per cycle. However, due to constraints in AMD compilers and FPGA architectures, the total transfer size must be divisible by a power of two. Padding is therefore introduced to align the data size accordingly and enable more efficient transfers.
In the code presented in Listings 2 and 3, where $J=190$, transferring all elements along the second dimension (iterated by loop $j$) of array $B$ onto the chip is constrained by the available memory bandwidth. Without padding, the maximum transfer rate is $64$ bits per cycle, as $190\times 32$ is divisible by $64$ but not by $128$. However, by introducing padding with $P=2$ and adjusting $J$ to $192$ ($190+2$), the data alignment enables a significantly higher transfer rate of $512$ bits per cycle, as $192\times 32$ is now perfectly divisible by $512$. This optimization maximizes memory bandwidth utilization, reducing transfer latency and improving overall data throughput.
@@ 12 A[i] += 1;
Padding can also be used for computation, if we take a similar method than Sisyphus (Pouget et al., 2025b) which tiled and unroll the intra-tile loops which correspond to the transformation from Listing 2 to Listing 3. We will have an unroll factor which correspond to $I1\times J1\times K1$ but we want to have $I1$ divide $I$, $J1$ divide $J$ and $K1$ divide K because we do not want to use extra resource to only compute a partial tile which would correspond to the execution of the last tile which is not complete. For example if we have a trip count of $190$ and we use an unroll factor of $8$ then we will execute $184\times 8$ a full tile and then 6 iterations for the partial tile.
If the unroll factor is restricted to be a divisor of the loop trip count, the design space becomes significantly limited. Padding can be used to adjust the trip count, thereby expanding the set of valid unroll factors to better match available hardware resources. As shown in Listing 1, a loop with a trip count of 190 permits only a limited set of unroll factors: $UF\in\{1,2,5,10,19,38,95,190\}.$ However, by padding the loop to a trip count of 192, the space of legal unroll factors becomes: $UF\in\{1,2,3,4,6,8,12,16,24,32,48,64,96,192\}.$ This increased flexibility allows finer control over unrolling and resource utilization, enabling more efficient hardware implementations and better exploitation of computational parallelism.
@@ 4 C[i][j] += A[i][k] * B[k][j];
Listing 2: Baseline Implementation of Matrix Multiplication in C. This code depicts a naive triple-nested loop structure used for matrix-matrix multiplication, serving as the unoptimized reference for further transformations.
@@ 10 C[i0*I1+i1][j0*J1+j1] += A[i0*I1+i1][k0*K1+k1]*B[k0*K1+k1][j0*J1+j1];
Listing 3: Matrix Multiplication with Loop Tiling and Fully Unrolled Intra-Tile Computation. This implementation showcases a performance-optimized matrix multiplication using loop tiling to enhance data locality and fully unrolled intra-tile loops to expose fine-grained parallelism.
@@ 2.3. Challenges
A common oversimplification in many frameworks is the assumption that all data resides on-chip e.g., (Basalama and Cong, 2025; Ye et al., 2022). While this simplifies memory access patterns in design space exploration, it often results in low Quality of Results (QoR) when deployed on real hardware. This assumption prevents frameworks from incorporating crucial tiling strategies, which are essential for optimizing data locality, reducing memory access overhead, and ensuring efficient off-chip communication. When off-chip memory management is treated as a separate optimization step, decisions made in earlier phases may restrict the effectiveness of later memory optimizations, leading to suboptimal performance and resource utilization.
The limitations of current HLS optimization frameworks that rely on shared buffering are exemplified through the 3mm kernel. It consists of two independent matrix multiplications, whose outputs serve as inputs for a final matrix multiplication. The 3mm kernel computes: G = (A × B) × (C × D) where intermediate matrices E and F store the results of the two first multiplications.
As illustrated in Listing 4, each matrix multiplication consists of deeply nested loops that are well-suited to parallelization techniques such as loop unrolling, pipelining, and computation-communication overlap. However, conventional shared buffering strategies often fall short in effectively leveraging concurrency, primarily due to their limited exploitation of dataflow principles and insufficient overlap between computation and communication.
@@
Allo (Chen et al., 2024) aims to reduce development time and enhance design quality through a composable programming model for hardware accelerator design. By decoupling algorithm specification from hardware customization and enabling modular schedule composition, Allo supports holistic dataflow optimization and verifiable transformations for large-scale, multi-kernel designs. However, it still requires significant manual intervention from an expert, limiting its usability for non-specialists.
@@ 15 G[i][j] += E[i][k]*F[k][j];}// S5
Listing 4: Reference Implementation of the 3mm Kernel from PolyBench. The 3mm kernel computes: G = (A × B) × (C × D) where intermediate matrices E and F store the results of the two first multiplications.
1for (i = 0; i < 180; i++) // Task 0
2 for (j = 0; j < 190; j++)
3 E[i][j] = 0.0; // S0
4for (i = 0; i < 180; i++) // Task 1
5 for (j = 0; j < 190; j++)
6 for (k = 0; k < 200; ++k)
7 E[i][j] += A[i][k]*B[k][j];}// S1
8for (i = 0; i < 190; i++) // Task 2
9 for (j = 0; j < 210; j++)
10 F[i][j] = 0.0; // S2
11for (i = 0; i < 190; i++) // Task 3
12 for (j = 0; j < 210; j++)
13 for (k = 0; k < 220; ++k)
14 F[i][j] += C[i][k]*D[k][j];}// S3
15for (i = 0; i < 180; i++) // Task 4
16 for (j = 0; j < 210; j++)
17 G[i][j] = 0.0; // S4
18for (i = 0; i < 180; i++) // Task 5
19 for (j = 0; j < 210; j++)
20 for (k = 0; k < 190; ++k)
21 G[i][j] += E[i][k]*F[k][j];}// S5
Listing 5: Direct Transformation of Listing 4 Where Each Loop Body Becomes a Separate Task
@@ 2.4. Overview of Prometheus
On-Chip Memory Available Total memory usage by buffered tiles (considering reuse and buffering) must be within BRAM capacity.
@@
Figure 2 illustrates the workflow of Prometheus. The process begins with an input C++ code, from which we extract an intermediate representation containing dependencies, trip counts, schedules, and other relevant information. This extraction is performed using PoCC (PoCC, [n. d.]), which provides all necessary details for analysis. Next, we construct the task-flow graph based on this extracted information and formulate the corresponding Non-Linear Programming (NLP) problem. Solving this NLP problem determines the theoretical optimal parameters for the design space. Once the NLP solver provides these parameters, we have all the required information to proceed. At this stage, we automatically generate the HLS-C++ code along with all the necessary files to produce the FPGA bitstream, ensuring an efficient and fully automated compilation process.
@@
The pseudo-code in Listing 6 demonstrates how Prometheus processes the 3mm kernel (Listing 4). Load and store operations manage data transfers to and from off-chip memory, while send and receive operations handle inter-task communication using FIFO.
Each taski in the pseudo-code corresponds to the fully unrolled computation of an intra-tile for statement Si in Listing 4. An example of such a task is illustrated in Listing 7.
@@
Prometheus enhances computation efficiency by fusing statements that produce the same outputs (e.g., in Listing 4), and it automatically formulates an NLP problem to determine the theoretically optimal parameters, such as loop schedules, array bit widths (e.g., 512 bits), tile sizes, reuse buffer sizes, transfer locations, and padding. The framework dynamically adjusts bit widths for arrays like $F$, $D$, and $G$ to achieve a better balance between parallelism and resource usage.
@@
Listing 6: Transformed and Fused 3mm Kernel Pseudocode Generated by Prometheus. This code highlights the result of Prometheus optimizations, including fused task generation, tiled loops, memory buffer management, and pipelined execution with dataflow pragmas.